OpenStudy (anonymous):

What is the factored form of the expression? 3x2 + 8x – 16

OpenStudy (mathmate):

3x^2+8x-16 =x(3x-4)+12x-16 =x(3x-4) +4(3x-4) =(x+4)(3x-4)

OpenStudy (anonymous):

thank you

OpenStudy (mathmate):

You're welcome!